-9(x-3)^2=-7, can someone please help me I dont understand how to do this one. Thank you in advance!
Answer by edjones(8007)

-9(x-3)^2=-7
-9/-9(x-3)^2=-7/-9
(x-3)^2=7/9
x-3=+-sqrt(7/9) Sqrt of each side.
x-3=+-sqrt(7)/3
x=3+-sqrt(7)/3
.
Ed